What is the work done by gravity when a wooden block of mass 2 kg slides 3 m down a slope inclined at 30°?

A. 3 J

B. 29.43 J

C. 6 J

D. 58.86 J


1
Expert's answer
2022-02-01T11:57:30-0500

m=2kg

Angle"(\theta)=30°"

h=3m

Now We know that


"V=\sqrt{2gh}=\sqrt{2\times9.8\times3}=7.66m"

Work done by gravity

"W=mghsin\theta"


"W=2\times9.8\times3\times sin30°=29.43J"

Option (b) is correct option